Advance Study Assignment: Reaction Stoichiometry: The Mole Ratio Concept 1. What is the molar mass of Fe(NH(SO 6 H,0? 2. What mass of Fe(NHJSOJ 6 H,o is required to prepare 50.0 mL of a 1.50 x 10 M solution? 3. What volume (mL) of 6.0 M H,SO, is required to prepare 50.0 ml. of a solution which is 0.25 M in H,SO,? How many drops is this? (Assume 1 mL.-20. drops.) 4. In the reaction between OCI and Crt in acidic solution, the following data is obtained. Plot the absorbance of the solution (points only) as a function of the OCT/Cr mole ratio. Perform a linear regression on each of the two linear segments and obtain the equation for cach line. Solve the equations simultancously to determine the mole ratio at which they intersect. Use this mole ratio to balance the equation for the reaction. OCI /CP mole ratio Absorbance at 430 nm 0.0 0.5 1.0 1.5 2.0 2.5 3.0 0.044 0.168 0.288 0.416 0.419 0.422 0.425 What is the mole ratio of OCI to Cr in this reaction? 4. Using this mole ratio, balance the following equation. Remember, coefficients are whole numbers.
Answer #1

Q1

MW = total moles in specie

MW = Fe + 2x(NH4) + 2x(SO4) + 6x(H2O)

MW = 55.85 + 2*18 + 2*96.0626 + 6*18

MW = 391.9752 g/mol

Q2.

mol = M*V

M = 1.5*10^-2

V = 50*!0^-3 L

mol = M*V = (1.5*10^-2)(50*10^-3) = 0.00075

mass = mol*MW = 0.00075*391.9752 = 0.2939814 g of salt

Q3

20 drops = 1 mL

x drops = 50 mL

x = 50*20 = 1000 drops